OpenStudy (anonymous):

Athena uses her credit card to buy some clothes for $738.14. She can pay off up to $265 per month. The card has an annual rate of 17.9% compounded monthly. How much total interest will she pay?

OpenStudy (anonymous):

$10.27 $21.61 $26.04 $11.01

OpenStudy (anonymous):

@satellite73 @jim_thompson5910 @thomaster

OpenStudy (anonymous):

@Compassionate @AccessDenied @whpalmer4

ganeshie8 (ganeshie8):

Balance after 1 month : \(738.14\left(1+ \frac{0.179}{12} \right) - \color{red}{265} = ?\)

OpenStudy (anonymous):

i got 484.15

ganeshie8 (ganeshie8):

yes thats the balancce after 1 month

ganeshie8 (ganeshie8):

calculate the balance after 2 months

ganeshie8 (ganeshie8):

Balance after 1 month : \(738.14\left(1+ \frac{0.179}{12} \right) - \color{red}{265} = 484.15\) Balance after 2 months : \(484.15\left(1+ \frac{0.179}{12} \right) - \color{red}{265} = ?\)

OpenStudy (anonymous):

i got 21.60 as the final answer, is that right?

ganeshie8 (ganeshie8):

21.6 is \(\large \color{red}{\checkmark}\)